Project

General

Profile

Actions

Meeting 2014-09-03

Date: 3rd September 2014
Time 15:00 BST
Attendance:
Apologies: A. Dobbs

Agenda and Minutes

0. Actions

  • Email AB about what spare waveguide material is available - GB
    DONE
    Spare waveguide material at Imperial. 3 or 4 weeks.
  • Asymmetric results between trackers: Kalman to extrapolate back to tracker reference surface - C. Hunt, ES
    Not available for comment.
  • Create MICE note of channel mappings end-to-end through hardware, electronics and software - DA
    In progress.

AFE/cassette channel mapping: MICE note 447.

  • DAQ and controls requirements for schedule - DA CM talking to DA
    Done
  • Create comparison plots for channel mapping data - CH
    Unavailable

1. Software: AD, ES, CH, ST

  • Schedule - AD See MAUS Schedule.
  • MAUS version 0.9.0 released with lots of scifi updates

No update

1.1 Documentation - AD, ES, CH

  • Base documentation - AD
  • Digitisation - CH, ES
  • Pattern Recognition - AD
  • Track fit - ES
  • MC - CH
  • Geometry, CDB - CH
  • Noise - CH

No update

1.2 Geometry, Calibrations, Configurations, CDB - CH

  • Helium windows not in geometry - CH
  • Get CMM geometry data in to CDB - CH
  • Translate CDB geometry xml into MiceModules format Now able to pull down from CDB but still crashing when try to run sim with it - AD, C. Heidt

1.3 Monte Carlo

  • MC Noise - CH
    • Look at large cpu load
  • ADC smearing needs fixing - CH

1.4 Pattern Recognition - AD

  • Find a way increasing error for low pt tracks and implement

1.5 Online Displays - MU

1.6 Performance

  • Spacepoint efficiency study - CH, AD
  • Bias difference between trackers to be investigate - C. Hunt, ES 99% solved, need to get Kalman to extrapolate back to tracker reference surface
  • Low pt performance - AD
  • Recon emittance vs MC truth emittance study - C. Hunt
  • Spacepoint efficiency study - C. Hunt

1.7 Issues

  • Final track fit - ES
    • Deal with -ve charge
  • Possible bug found in plane mapping

2. DAQ/Calibration/Controls - DA, EO

  • DAQ system test in Nov 14?
  • DA to give his requirements for the Hall and the tracker lab, in particular with respect to the Integration Schedule (MU to email)
    DA coming to U.K. in first two weeks of September

    Plan for QA it September doc
    Plan for integration with beam doc

One of the cryos has been cabled up so we have one board readout electronics all cabled up. No problems yet. Running, initialises and reads out compressed data Pedestal data taken (as cryos not cold) 1 run taken with bayonet out and one in (shows load above pedestal) visible in ADC distribution. Boards show no major problems with chips etc. Rest of boards will now be checked. All checks that can be done while warm will be done. VLSB and cables all cycled through.

Vacuum pumps brand new but all broken. Will be serviced by Craig in due course.

2.1 MiniDAQ - DA

Being used.

2.2 Channel Mappings

  • Check our channel mapping from end to end and make sure correct - DA, AD, ES, CH
  • Dig up the old documentation for the electronics channel mapping - DA

came from concerns that channels were swapped after cosmics. We think that it is not a problem.

ACTION: David A. to email Adam.
ACTION; AB Find D0 note and make into MICE note.

2.3 Controls

  • Create controls system - DA

2.4 Hall Probes - MU

We already have a working system but we will improve it and make it ore user friendly. But as

3. LED: EO

  • EO currently busy with thesis, back on LED in Sept - Oct
  • Finished second pulser box - CM
  • Polish fibre ends - CM
  • Make documentation mapping LEDs to areas of tracker illuminated - EO
  • Provide documnetation to Maurizio about whether it would be helpful to fit LED into TOF to help our timing - EO

Craig has the bare bones of the system and Ed will assemble the spares when he has finshed writing and check everything in more detail.

4. Tracker lab - AB, GB, CM

  • Set up cryostat test system - CM
  • Turbo overhaul?
  • Helium line in from R9 now in lab and working well

Tidy and in use.

5. Installation

  • ISIS long shut down: 25th Aug 2014 - 8th Feb 2015
  • Possible helium leak on the downstream solenoid
    Observed at the Wang and not an issue. There was just some confusion with the data but it is resolved and not an issue.
  • Expecting to move downstream SS to Hall on Monday
    done.
  • Tracker 2 log to be produced by GB with updated numbering
  • Schedule
    • Current schedule at Integration Schedule - timetable still subject to change (constraints from AFC, etc)
    • Add resource loading? - CM
  • Finish installation manual - GB, MU Done, stick final version on wiki
  • External waveguides
    • September to be begin QA

Work all set to begin. New software almost complete. Initail calibration tests complete. Run plan complete. Real work to begin in nest two weeks and to take 4 weeks data taking followed by a few weeks analysis.
Efficiency numbers to be added to electronics etc efficiency numbers from cosmics running and included in the SW. Any completely broken (or damaged to outside tolerance to be replaced).

  • Load of fibre and 5 complete connectors random received from AB, not too sure what to do with them. GB to email AB about what is available to determine if we need to order items in.

6. Services in the MICE hall: CM

  • CM discussing with AN and JG about chilled water and compressors
  • Helium lines

No update.

7. Step IV Analysis Readiness - MU

  • Tracker alignment - MU

MC is still being developed by C Heidt to include the helium window. But work is progressing. Upon completion the simulation and analysis, following the plans discussed in previous meetings, will be done.
Updates on a similar timescale to the Waveguide QA as Melissa will undertake work simultaneously.

8. Risks

  • Daresbury electronics: vacuum, tracker interlocks
  • Reforming external waveguides (could lead to breakage)
  • Leak from cryostat cassettes
    3/4 tested and tight. AB to test the 4th on 3rd sept.
  • Northside PRY (not us, but knock effect to us)
    Can we add cryostats etc early..? decision to be made.
  • Not yet able to test against multiple tracks per trigger with MC

9. AoB and summary of actions.

Phone Details

  • The four digit company id: 4673 (callers from outside the UK)
  • The pin number: 419357

Phone numbers:

UK (Imperial College --preferred) 0845 545 4460 English
Austria 0820 4000 1503 English
0820 4000 15 02 German
Belgium 070 35 99 45 English
070 35 99 94 Dutch
070 35 98 66 French
Canada 1 712 432 2832 English
1 712 432 2844 French
Germany 01803 001 178 English
01570 248 8303 English (mobile)
01803 001 177 German
01570 248 8101 German (mobile)
Ireland 0818 270 007 English
Italy 848 390 166 English
848 391 819 Italian
Netherlands 0870 001 901 English
0870 001 909 Dutch
Poland 0801 003 543 English
0801 003 533 Polish
South Africa 087 550 0375 English
Spain 902 885 318 English
902 881 200 Spanish
Sweden 0939 2066 400 English
0939 2066 300 Swedish
Switzerland 0848 560 190 English
0848 560 378 French
0848 560 347 German
0848 560 385 Italian
United Kingdom (PowWowNow) 0844 4 73 73 73 English
United States 1 415 363 0833 English

Updated by Dobbs, Adam about 9 years ago · 4 revisions